OpenStudy (anonymous):

Find all z in C which satisfy z conjugate = z^2

OpenStudy (anonymous):

Note C=Complex (x+iy)

OpenStudy (saifoo.khan):

@sat http://openstudy.com/#/updates/4ed58b7be4b0bcd98ca6ebe8

OpenStudy (anonymous):

0 works 1 works i don't think anything else works. \[(a+bi)^2=a^2-b^2+(2ab)i\] so this would mean you would have to solve \[a=a^2-b^2\] and \[b=2ab\]

OpenStudy (anonymous):

sorry it is \[\overline{z }= z^2\] so it would be \[a=a^2-b^2\]and \[-b=2ab\] but the outcome is still the same

OpenStudy (anonymous):

i got a=-1/2 and b=\[\sqrt{3}/2\]